INAF queue observing, 2014 Oct 27-28

Observer: E. Sani, R. Carini
Telescope Operator: David Gonzalez-Huerta
Support Astronomer: Michelle Edwards
Instrument: MODS

Summary

We start observations at 1:45 UT<br />
At the beginning of observations we have thin cirrus above the orison (NO direction), low wind ans humidity at 26%.
N.B. The seeing from the GCS is overestimated by 10% because the MODS guider is out of focus.
During night humidity is <10%, no wind, clear sky, variable seeing: 0.8 - 1.1 in red

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard Feige110 dual grating + red grating - airmass 1.4. Default 5" slit
Fusi Pecci Q B517 dual grating 45m COMPLETED
Fusi Pecci R H2 dual grating 45m COMPLETED
Saracco F XLLSM1 red grating 2.30 h parallel mode failed
Dotto KC46 V- ,R -Bessel 1 h  
Clementini IZw 18 V- I- Bessel 10m  

Details

01:40 ready to operate with MODS
  • we slew to the standard star Feige110

01:45 acqMODS feige110.acq mods1r.20141028.0001-0002

2:05 R. Pogge try to operate direcly with the MODS computer to try to figure out the ds9 problem. All the ds9 processes are killed and we restart ds9.
  • modsAlign -r mods1r.20141028.0002.fits
  • execMODS feige110.obs mods1r.20141028.0003
  • we can not use modsDisp because it has no more writing permissions in newdata. We have to display each image manually from newdata and from obs2 because there is some conflict in obs3.
  • mods1r.0008 is corrupted, we run again a red-grating-only spectrum
  • mods1b.003-005 mods1r.0005-007 dual grating
  • mods1r.007, 009, 0010 red grating only

2:40 we begin Fusi Pecci Q Program
  • acqMods B517.acq mods1r.0011-12
  • modsAlign -l mods1r.0011.fit mods1r.0012.fits mods1r.0013-0014

3.03 execMODS B517_ls_06.obs =mods1r.0015-0017==mods1b.0006-0008 =
  • we begin the scientific acquisition with FWHM mean=1.4
  • during the observation FWHM is diminished to 0.9

03:58 acqMods H2_ls.acq = mods1r.0018-0020=
  • abort the preset to unwrap the rotator

04:03 acqMods H2_ls.acq modsAlign -l mods1r.0021.fit mods1r.0022.fits mods1r.0023

04:20 execMODS H2_ls_06.obs mods1r.0024-26, mods1b.0009-0011
  • at the beginning of the integration the average FWHM on the GCS is 0.85

5:15 we begin the parallel mode for Saracco F program *acqMODS mods.5.xllsm1.acq = mods1r.0027=

5.19 dofpia, /X2, /first, /RedOnly on the MODS pointing
  • error floating divide by 0
  • the problem is that the procedure does not find 7 pupils and this happens because the Telescope box in the OB execution panel is authomatically uncheked when dofpia runs.
  • We try committing clearing the optics, committing global offsets to the primary mirrors as computed by Jhon, we gave coma corrections by hand to allow dofpia to see the pupil, but the problem persists because it is actually due to some bug. The bug happened possibly because we ran dofpia without the guiding.
  • Thus we do apply a new approach: reset everithing using both cameras and running biases with Telescope and Guiding boxes cheked. Run dofpia. Uncheck the guiding.
  • John call back and says probably he applied the coma and commit new corrections to David.
  • We thus don't actually where and when the bud originated.

05:40 We are loosing time, so we start with modsAlign mods.5.xllsm1.mms mods1r.0027 mods1r.0028 mods1r.0029-0037

06:05 dofpia, /X2, /RedOnly converged
  • Run XLSSm1_I_1_180s.ob
  • images are elongated. This seems to arise from a wrong center for the rotator, indeed there is no co-pointing in pseudo parallel mode. Anyhow the central 4'x4' in chip 2 (i.e. the corrected FoV of MODS) are fine

06:12 execMODS mods.5.xllsm1_1.obs
  • due to a rgo command in the script, we get errors. We do rerun the execmods, but since there is an updatepoing and and offsets was already done, we risk to send the scientific spectra out of the slit. We go back step by step and acquire through slit image to check the position.
  • While doing this we see that while MODS is doing nothing, the LBC try to send guinding correction and the through slit images we are collecting are blurry.
  • This can affect the scientific frames in MODS so we decide to DISCONNECT LBT and ABORT PSEUDO-PARALLEL MODE

06:37 execMODS mods.5.xllsm1_1.obs mods1r.0038-0041
  • the average seeing at the beginning is 1.1 on the GCS

07:48 execMODS mods.5.xllsm1_2.obs mods1r.42-0045
  • the average seeing at the beginning is 0.9 on the GCS
  • images mods1r.0043, mods1r.0045 appears corrupted in /newdata but is correct on the MODS computer.
  • This happens when the ds9 opened from /newdata unespectedly crashes. We recover the images: /newdata/mods1r.20141028.0043r.fits, mods1r.20141028.0045r.fits

09:00 execMODS mods.5.xllsm1_3.obs mods1r.46-0047
  • only half OB because we are reaching the airmass limit
  • mods1r.0047 is saved in mods1r.20141028.0047r2.fits

09:40 We ask the TO to switch to LBC binocular

10.02 We are authorized to use LBC
  • Slew on Dotto_DDT KC46
  • According to the ephemeridis we run collimation and copointing on a nearby star: WT10-271

10:29 Run 2014KC46.ob

11:50 Slew to Clementini_10
  • dofpia, /X2 on the copointing field
  • dofpia converged but then the primary morrors reach their limits: too many forces applied. We have to clean the optics and re-run dofpia
  • lbcrangebal B=121315 R=121309
  • preset canceled
  • rerun lbcrangebal

12:17 Run IZw18_3.ob

12:35 We star skyflats
  • Run g-Sloan I-Bessel

13:22 The TO closes the Dome
  • we Run MODS calib for Fusi_Pecci_17 program
  • grpixflat.cal mods1b.0012-0021 mods1r.0048-0052 SlitLess flats, DUAL grating
  • slitflats.cal mods1b.0022-27 mods1r.0053-0058 1.2" Slit flats, DUAL grating
  • frlamps.cal mods1b.0028-0030 mods1r.0059-61 Arc, DUAL grating 0.6" slit
  • frlamps.cal mods1b.0031-0032 mods1r.0062- Arc, DUAL grating 1.2" slit

14:42 execMODS modsSleep.pro

%USERSIG{RobertaCarini - 2014-10-28}%

Comments


 
Topic revision: r11 - 18 Dec 2014, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback